OpenStudy (anonymous):

2(2c)-6(c-1)=-5(c+3)+11 solve symbolically < or>

OpenStudy (anonymous):

Facts you may need: \(a(b+c)=ab+ac\) \(\{+,\times\}\) has the set of inverse operations \(\{-,\div\}\).

OpenStudy (anonymous):

i broke it down to 3c =-10 but dont know if its right or what to do

OpenStudy (anonymous):

I think you made an error somewhere. Check your calculations.

OpenStudy (anonymous):

ok let me check

OpenStudy (anonymous):

cant figure it out

OpenStudy (anonymous):

One moment

OpenStudy (anonymous):

Sorry! I had made an error myself. You were correct. All you do now is divide by \(3\). I am so sorry.